OpenStudy (anonymous):

solve the equation 4=√p-2

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

i think is 36

OpenStudy (anonymous):

you want to be friends Chiko i fanned you

OpenStudy (yanasidlinskiy):

@hatiekate1232314 welcome to OpenStudy although this isn't a place to say that.

OpenStudy (anonymous):

ummm..ok...

OpenStudy (anonymous):

maybe it's 18, let me see

OpenStudy (anonymous):

yeah sooo...was i right???

OpenStudy (yanasidlinskiy):

\[4= \sqrt{p-2}\] Like that?

OpenStudy (anonymous):

yup i think you got it

OpenStudy (yanasidlinskiy):

Do NOT tell someone that they are right when they're wrong. That is plain rude.

OpenStudy (anonymous):

yeah sooo...was i right???

OpenStudy (yanasidlinskiy):

Like I said, that is incorrect, Chiko.

OpenStudy (anonymous):

i was just about to say wait i think it was wrong i think its 10

OpenStudy (anonymous):

sorry i had to refresh my page...

OpenStudy (yanasidlinskiy):

Stop giving out direct and incorrect answers. It's fine, Chiko.

OpenStudy (anonymous):

ok

OpenStudy (yanasidlinskiy):

How can you get 10?!

OpenStudy (anonymous):

let me post the answers....

OpenStudy (anonymous):

wait he is right 36 4=p√−2 4+2=p√ 6=p√ 62=(p√)2 p=36

OpenStudy (anonymous):

a.√6 b.36 c.3 d.6

OpenStudy (anonymous):

thx @YanaSidlinskiy and @hatiekate1232314

OpenStudy (anonymous):

ya still here??

OpenStudy (anonymous):

4=p√−2 4+2=p√ 6=p√ 62=(p√)2 p=36

OpenStudy (anonymous):

ya i got it right..thx

OpenStudy (yanasidlinskiy):

Ask that in Messaging.

OpenStudy (anonymous):

lol ok...